A 24-foot-tall pole stands next to the highway. A 25-foot-long support wire stretches from the top of the pole to the ground . How far away from the base of the pole does the wire meet the ground
Answer by richwmiller(17219)

I find the question quite funny.
Of the three measurements (the height of the flag pole, the length of the wire and the distance from the pole), the easiest to measure by hand is the distance from the pole. Just pace the distance- you don't have to climb the pole nor the wire to measure the distance.
25^2=24^2+x^2
625-576=x^2
49=x^2
x=7